# Cartesian Nodal Lines and Magnetic Kramers Weyl Nodes in Spin-Split Antiferromagnets

### v1

@@ Abstract
When open boundary conditions are applied in the $(\bar{1}10)$ direction, $k_{2}$ and $k_{z}$ can be treated as parameters. The existence of topological boundary states can then be determined by analyzing the topological properties of the reduced one-dimensional Hamiltonian along a given $k_{1}$ line. For fixed values of $k_{2}$ and $k_{z}$, the Hamiltonian possesses both inversion symmetry and chiral symmetry. The inversion symmetry operator is given by $\mathcal{P}=\sigma_{-}$, which satisfies $\mathcal{P}\mathcal{H}_{\rm CNL}(k_{1},k_{2},k_{z})\mathcal{P}^{-1}=\mathcal{H}_{\rm CNL}(-k_{1},k_{2},k_{z})$. Similarly, the chiral symmetry operator is given by $\mathcal{S}(k_{2},k_{z})=-\sin\theta(k_{2},k_{z})\sigma_{+}+\cos\theta(k_{2},k_{z})\sigma_{z}$, which satisfies the anticommutation relation $\{\mathcal{S}(k_{2},k_{z}),\mathcal{H}_{\rm CNL}(k_{1},k_{2},k_{z})\}=0$. It is noteworthy that the inversion symmetry operator $\mathcal{P}$ does not depend on $k_{2}$ and $k_{z}$, while the chiral symmetry operator $\mathcal{S}(k_{2},k_{z})$ does. If we restore $k_{2}$ and $k_{z}$ as momentum components, $\mathcal{S}(k_{2},k_{z})$ can no longer be interpreted as a chiral symmetry operator because it depends on momentum. This contradicts the requirement for a chiral symmetry operator, which must be a constant unitary and Hermitian operator. When a unitary and Hermitian operator, which depends on partial momentum components and anticommutes with the Hamiltonian, exists, the Hamiltonian is said to possess subchiral symmetry, and the corresponding operator is referred to as a subchiral symmetry operator[94]. This concept turns out to be very useful for diagnosing the topology of the Hamiltonian on certain dimension-reduced closed manifolds and the properties of the associated topological boundary states.
@@
[…]
By a straightforward calculations, we find that
@@
provided that $k_{2}\neq\{0,\sqrt{2}\pi\}$ and $k_{z}\neq\{0,\pi\}$. It is readily seen that the winding number has two nontrivial values, ±$\pm 1$. Compared to the single value $\phi=\pi$, this suggests that the chiral symmetry can provide more information on the topological surface states. Indeed, since the chiral symmetry is a nonspatial symmetry, the zero-energy surface states also serve as the eigenstates of the chiral symmetry operator. Since the two eigenstates of the operator $\mathcal{S}(k_{2},k_{z})$ are straightforward to obtain, the spin texture of the surface states can readily be determined. Specially, since $\mathcal{S}(k_{2},k_{z})=-\sin\theta(k_{2},k_{z})\sigma_{+}+\cos\theta(k_{2},k_{z})\sigma_{z}$, it is straightforward to obtain its two eigenstates, which are
@@
where $\beta=+1$ ($-1$) refers to the left (right) surface. These results demonstrate that the spin polarizations of the topological surface flat bands are momentum-dependent. This behavior is fundamentally different from that of the topological surface flat bands in a nodal-line semimetal protected by chiral symmetry, where the spin polarizations are fixed and momentum-independent.
@@
$\mathcal{H}(\boldsymbol{k})$ $=$ $2\left[-t(\cos k_{x}+\cos k_{y})-t_{z}\cos k_{z}\right]\sigma_{0}+\lambda_{\rm so}\sin k_{x}\sigma_{x}+\lambda_{\rm so}\sin k_{y}\sigma_{y}+\lambda_{\rm so}\sin k_{z}\sigma_{z}.$ (S20)
@@
[…]
@@
When both $\lambda_{\rm M}$ and $\lambda_{\rm so}$ are finite, the band degeneracies in the band structure remain to be Weyl nodes, but the distributions of these Weyl nodes becomes a little more complex compared to that of a Kramers Weyl semimetal. First, because the $C_{4z}\mathcal{T}$ is preserved in the Hamiltonian, the four $C_{4z}\mathcal{T}$-invariant momentums, including $(0,0,0)$, $(0,0,\pi)$, $(\pi,\pi,0)$, $(\pi,\pi,\pi)$, are the locations of symmetry-enforced Weyl nodes. As these $C_{4z}\mathcal{T}$-invariant momentums are also TRIMs, we refer to these position-fixed Weyl nodes as magnetic Kramers Weyl nodes (MKWNs), highlighting their positions at TRIMs and the breaking of time-reversal symmetry.
To determine the distribution of potential additional Weyl nodes, we divide the analysis into two scenarios: one where $\lambda_{\rm M}>\lambda_{\rm so}$, and the other where $\lambda_{\rm M}<\lambda_{\rm so}$. Throughout, we consider $\lambda_{\rm M}$, $\lambda_{\rm so}$ and $\eta$ to be positive constants. Recall the conditions for the emergence of band degeneracies: (I) $\sin k_{x}(\lambda_{\rm so}-\lambda_{\rm M}\sin k_{z})=0$; (II) $\sin k_{y}(\lambda_{\rm so}+\lambda_{\rm M}\sin k_{z})=0$; (III) $\lambda_{\rm so}\sin k_{z}+\eta\lambda_{\rm M}(\cos k_{x}-\cos k_{y})=0$. We begin by examining the scenario where $\lambda_{\rm M}<\lambda_{\rm so}$, as it is simpler to analyze. For this case, condition (I) determines $k_{x}=0$ or $\pi$. Similarly, condition (II) determines $k_{y}=0$ or $\pi$. The results obtained under these two conditions indicate that the Weyl nodes must be located along the four high-symmetry $k_{z}$ lines at $(0,0,k_{z})$, $(0,\pi,k_{z})$, $(\pi,0,k_{z})$ and $(\pi,\pi,k_{z})$. On the two $C_{4z}\mathcal{T}$-invariant lines at $(0,0,k_{z})$ and $(\pi,\pi,k_{z})$, condition (III) reduces to $\lambda_{\rm so}\sin k_{z}=0$, which leads to $k_{z}=0$ and $\pi$, suggesting that only the four MKWNs appear on these two lines. On the line at $(0,\pi,k_{z})$, condition (III) simplifies to: $\lambda_{\rm so}\sin k_{z}+2\eta\lambda_{\rm M}=0$, which has solutions only if $\lambda_{\rm so}>2\eta\lambda_{\rm M}$. When this condition is fulfilled, two Weyl nodes emerge at $(0,\pi,-k_{0})$ and $(0,\pi,-\pi+k_{0})$, where $k_{0}=\arcsin 2\eta\lambda_{\rm M}/\lambda_{\rm so}$. Similarly, on the line at $(\pi,0,k_{z})$, two additional Weyl nodes emerge at $(\pi,0,k_{0})$ and $(\pi,0,\pi-k_{0})$, provided that $\lambda_{\rm so}>2\eta\lambda_{\rm M}$.
@@
It is easy to see that besides the Weyl nodes discussed in the first scenario, potential additional Weyl nodes may emerge along the following lines:
@@
(1) On the line at $(k_{x},0,\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}})$: Condition (III) reduce to
@@
Figure S1: Schematic of the magnetic order. White and blue spheres refer to nonmagnetic and magnetic atoms, respectively. The red arrows represent the orientations of the localized magnetic moments.
@@
In this section, we present a specific magnetic configuration as an example to demonstrate a possible realization of the tight-binding Hamiltonian under consideration. The distribution of the local magnetic moments is presented in Fig. S1. The white spheres represent nonmagnetic atoms, which supply itinerant electrons, while the blue spheres denote magnetic atoms that create a magnetic-order background for these electrons. When the hopping paths of the itinerant electrons intersect with the localized magnetic moments, the moments generate an effective spin-dependent potential. This potential, in turn, gives rise to spin-dependent hopping amplitudes. Based on the magnetic configuration illustrated in Fig. S1, the tight-binding Hamiltonian describing the itinerant electrons (assuming a single orbital degree of freedom for these electrons) is given by
$H$ $=$ $-\sum_{\braket{i,j},\sigma}t_{ij}c_{i,\sigma}^{\dagger}c_{j,\sigma}-\mu\sum_{i,\sigma}c_{i,\sigma}^{\dagger}c_{i,\sigma}+i\lambda_{\rm so}\sum_{\braket{i,j},\sigma,\sigma^{{}^{\prime}}}\boldsymbol{d}_{ij}\cdot c_{i,\sigma}^{\dagger}\boldsymbol{\sigma}_{\sigma\sigma^{{}^{\prime}}}c_{j,\sigma^{{}^{\prime}}}$ (S35)
$+\lambda_{\rm M}\eta\sum_{\braket{i,j},\sigma\sigma^{{}^{\prime}}}\boldsymbol{S}_{ij}\cdot c_{i,\sigma}\boldsymbol{\sigma}_{\sigma,\sigma^{{}^{\prime}}}c_{j,\sigma^{{}^{\prime}}}+\lambda_{\rm M}\sum_{\braket{\braket{i,j}},\sigma\sigma^{{}^{\prime}}}\boldsymbol{S}_{ij}\cdot c_{i,\sigma}\boldsymbol{\sigma}_{\sigma\sigma^{{}^{\prime}}}c_{j,\sigma^{{}^{\prime}}}+h.c.$
$=$ $\sum_{\boldsymbol{k}}\Psi_{\boldsymbol{k}}^{\dagger}\mathcal{H}(\boldsymbol{k})\Psi_{\boldsymbol{k}}.$ (S36)
@@
Here, $c_{i,\sigma}(c_{i,\sigma}^{\dagger})$ represents the annihilation (creation) operator for an electron with spin $\sigma$ at site $i$. The notation $\braket{i,j}$ indicates nearest-neighbor hopping between sites $i$ and $j$, $\braket{\braket{i,j}}$ indicates next-nearest-neighbor hopping, and the unit vector $\boldsymbol{d}_{ij}$ points along the bond direction from site $j$ to site $i$. The parameter $t_{ij}$ refers to the hopping amplitude between two nearest-neighbor sites, $\mu$ is the chemical potential, $\lambda_{\rm so}$ quantifies the strength of SOC, and $\lambda_{\rm M}$ characterizes the difference in hopping amplitude for opposite spins that is induced by the background magnetic moments. Performing a Fourier transformation to the momentum space and choosing the basis as $\Psi_{\boldsymbol{k}}^{\dagger}=(c_{A,\uparrow,\boldsymbol{k}}^{\dagger},c_{A,\downarrow,\boldsymbol{k}}^{\dagger},c_{B,\uparrow,\boldsymbol{k}}^{\dagger},c_{B,\downarrow,\boldsymbol{k}}^{\dagger})$, where $A$ and $B$ label two distinct sublattices within a unit cell, we obtain the momentum-space Hamiltonian, which reads
$\mathcal{H}(\boldsymbol{k})$ $=$ $-\left[t(\cos k_{x}+\cos k_{y})+t_{z}\cos k_{z}\right]\tau_{x}+\lambda_{\rm so}(\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y}+\sin k_{z}\sigma_{z})\tau_{x}$ (S37)
$+\lambda_{\rm M}\eta(\cos k_{x}-\cos k_{y})\sigma_{z}\tau_{x}+\lambda_{\rm M}\sin k_{z}(-\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y}).$
Here, Pauli matrices $(\sigma_{x},\sigma_{y},\sigma_{z})$ and $(\tau_{x},\tau_{y},\tau_{z})$ act on the spin and sublattice degrees of freedom, respectively. For notational simplicity, hereafter the identity matrices in the spin and sublattice spaces are made implicit, and the lattice constants are set to unity. It is evident that $\tau_{x}$ serves as a good quantum number for the Hamiltonian since $[\tau_{x},\mathcal{H}(\boldsymbol{k})]=0$. Therefore, the Hamiltonian $\mathcal{H}(\boldsymbol{k})$ can be decomposed as a direct sum of two independent parts corresponding to the two eigenvalues of $\tau_{x}$. Specially, we have $\mathcal{H}(\boldsymbol{k})=\mathcal{H}_{\tau_{x}=1}(\boldsymbol{k})\oplus\mathcal{H}_{\tau_{x}=-1}(\boldsymbol{k})$, where
$\mathcal{H}_{\tau_{x}=\tau}(\boldsymbol{k})$ $=$ $-\tau\left[t(\cos k_{x}+\cos k_{y})+t_{z}\cos k_{z}\right]+\tau\lambda_{\rm so}(\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y}+\sin k_{z}\sigma_{z})$ (S38)
$+\tau\lambda_{\rm M}\eta(\cos k_{x}-\cos k_{y})\sigma_{z}+\lambda_{\rm M}\sin k_{z}(-\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y})$
with $\tau=\pm 1$. It is evident that the two-band Hamiltonian $\mathcal{H}_{\tau_{x}=1}(\boldsymbol{k})$ is identical to the tight-binding model investigated in the main text, and the other two-band Hamiltonian $\mathcal{H}_{\tau_{x}=-1}(\boldsymbol{k})$ exhibits properties similar to those of $\mathcal{H}_{\tau_{x}=1}(\boldsymbol{k})$.
Figure S2: The evolution of the Hall conductivity $\sigma_{xy}$ (in unit of $e^{2}/h$) with respect to $B_{z}$ for different hopping amplitudes [(a,c)] and chemical potential [(b,d)]. We set $t=t_{z}$, $\mu=0$ in [(a,c)] and $t=t_{z}=0$ in [(b,d)]. The strength of spin-orbit coupling $\lambda_{\rm so}$ is set to 00 in [(a-b)] and $0.2$ in [(c-d)], respectively. Common parameters are $\lambda_{\rm M}=1$ and $\eta=1$.
@@
In the main text, we have shown that, when $\lambda_{\rm so}=0$, a weak Zeeman field will change the CNLs to two nodal rings located at the two mirror-invariant planes at $k_{z}=0$ and $\pi$, leading to the occurrence of three-dimensional quantized anomalous Hall effects when the Fermi surface only contains the nodal rings. The underlying physics can be understood by the layer Chern number. Specially, from a dimension-reduction perspective, the two-dimensional layer Hamiltonian with a fixed $k_{z}$ describes a Chern insulator characterized by a Chern number of value $\mathcal{C}=1$ or $-1$ (depending on the direction of the Zeeman field) when $k_{z}\neq 0$ and $\pi$. Accordingly, each layer contributes to a Hall conductivity $\sigma_{xy}=\frac{e^{2}}{h}$. By summing the contribution from all layers, one obtains

### v3

@@ Abstract
When open boundary conditions are applied in the $(\bar{1}10)$ direction, $k_{2}$ and $k_{z}$ can be treated as parameters. The existence of topological boundary states can then be determined by analyzing the topological properties of the reduced one-dimensional Hamiltonian along a given $k_{1}$ line. For fixed values of $k_{2}$ and $k_{z}$, the Hamiltonian possesses both inversion symmetry and chiral symmetry. The inversion symmetry operator is given by $\mathcal{P}=\sigma_{-}$, which satisfies $\mathcal{P}\mathcal{H}_{\rm CNL}(k_{1},k_{2},k_{z})\mathcal{P}^{-1}=\mathcal{H}_{\rm CNL}(-k_{1},k_{2},k_{z})$. The chiral symmetry operator is exactly the $\mathcal{S}(k_{2},k_{z})=-\sin\theta(k_{2},k_{z})\sigma_{+}+\cos\theta(k_{2},k_{z})\sigma_{z}$, which satisfies the anticommutation relation $\{\mathcal{S}(k_{2},k_{z}),\mathcal{H}_{\rm CNL}(k_{1},k_{2},k_{z})\}=0$. It is noteworthy that the inversion symmetry operator $\mathcal{P}$ does not depend on $k_{2}$ and $k_{z}$, while the chiral symmetry operator $\mathcal{S}(k_{2},k_{z})$ does. If we restore $k_{2}$ and $k_{z}$ as momentum components, $\mathcal{S}(k_{2},k_{z})$ can no longer be interpreted as a chiral symmetry operator because it depends on momentum. This contradicts the requirement for a chiral symmetry operator, which must be a constant unitary and Hermitian operator. When a unitary and Hermitian operator, which depends on partial momentum components and anticommutes with the Hamiltonian, exists, the Hamiltonian is said to possess subchiral symmetry, and the corresponding operator is referred to as a subchiral symmetry operator[98]. This concept turns out to be very useful for diagnosing the topology of the Hamiltonian on certain dimension-reduced closed manifolds and the properties of the associated topological boundary states.
@@
$W^{(\bar{1}10)}(k_{2},k_{z})=\frac{1}{4\pi i}\int_{-\sqrt{2}\pi}^{\sqrt{2}\pi}dk_{1}{\rm Tr}\left[\mathcal{S}(k_{2},k_{z})\mathcal{H}^{-1}_{\rm CNL}(k_{1},k_{2},k_{z})\partial_{k_{1}}\mathcal{H}_{\rm CNL}(k_{1},k_{2},k_{z})\right].$ (S11)
By straightforward calculations, we find that
@@
provided that $k_{2}\neq\{0,\sqrt{2}\pi\}$ and $k_{z}\neq\{0,\pi\}$. It is readily seen that the winding number has two nontrivial values, $\pm 1$. Compared to the single value $\phi=\pi$, this suggests that the chiral symmetry can provide more information on the topological surface states. Indeed, since the chiral symmetry is a nonspatial symmetry, the zero-energy surface states also serve as the eigenstates of the chiral symmetry operator. Since the two eigenstates of the operator $\mathcal{S}(k_{2},k_{z})$ are straightforward to obtain, the spin texture of the surface states can readily be determined. Specifically, since $\mathcal{S}(k_{2},k_{z})=-\sin\theta(k_{2},k_{z})\sigma_{+}+\cos\theta(k_{2},k_{z})\sigma_{z}$, it is straightforward to obtain its two eigenstates, which are
@@
where $\beta=+1$ (1-1) refers to the left (right) surface. These results demonstrate that the spin polarizations of the topological surface flat bands are momentum-dependent. This behavior is fundamentally different from that of the topological surface flat bands in a nodal-line semimetal protected by chiral symmetry, where the spin polarizations are fixed and momentum-independent.
@@
$\mathcal{H}(\boldsymbol{k})$ $=$ $\left[-t(\cos k_{x}+\cos k_{y})-t_{z}\cos k_{z}\right]\sigma_{0}+\lambda_{\rm so}\sin k_{x}\sigma_{x}+\lambda_{\rm so}\sin k_{y}\sigma_{y}+\lambda_{\rm so}\sin k_{z}\sigma_{z}.$ (S21)
@@
$\mathcal{C}_{(n_{1},n_{2},n_{3})\pi}=\frac{1}{2\pi}\oint\boldsymbol{\Omega}_{(n_{1},n_{2},n_{3})\pi}^{(c)}\cdot d\boldsymbol{S}.$ (S23)
@@
When both $\lambda_{\rm M}$ and $\lambda_{\rm so}$ are finite, the band degeneracies in the band structure remain to be Weyl nodes, but the distributions of these Weyl nodes becomes a little more complex compared to that of a Kramers Weyl semimetal. First, because the $C_{4z}\mathcal{T}$ is preserved in the Hamiltonian, the four $C_{4z}\mathcal{T}$-invariant momentums, including $(0,0,0)$, $(0,0,\pi)$, $(\pi,\pi,0)$, and $(\pi,\pi,\pi)$, are the locations of symmetry-enforced Weyl nodes. As these $C_{4z}\mathcal{T}$-invariant momentums are also TRIMs, we refer to these position-fixed Weyl nodes as magnetic Kramers Weyl nodes (MKWNs), highlighting their positions at TRIMs and the breaking of time-reversal symmetry.
To determine the distribution of potential additional Weyl nodes, we divide the analysis into two scenarios: one where $\lambda_{\rm M}>\lambda_{\rm so}$, and the other where $\lambda_{\rm M}<\lambda_{\rm so}$. Throughout, we consider $\lambda_{\rm M}$, $\lambda_{\rm so}$ and $\eta$ to be positive constants. Recall the conditions for the emergence of band degeneracies: (I) $\sin k_{x}(\lambda_{\rm so}-\lambda_{\rm M}\sin k_{z})=0$; (II) $\sin k_{y}(\lambda_{\rm so}+\lambda_{\rm M}\sin k_{z})=0$; (III) $\lambda_{\rm so}\sin k_{z}+\eta\lambda_{\rm M}(\cos k_{x}-\cos k_{y})=0$. We begin by examining the scenario where $\lambda_{\rm M}<\lambda_{\rm so}$, as it is simpler to analyze. For this case, condition (I) determines $k_{x}=0$ or $\pi$. Similarly, condition (II) determines $k_{y}=0$ or $\pi$. The results obtained under these two conditions indicate that the Weyl nodes must be located along the four high-symmetry $k_{z}$ lines at $(0,0,k_{z})$, $(0,\pi,k_{z})$, $(\pi,0,k_{z})$ and $(\pi,\pi,k_{z})$. On the two $C_{4z}\mathcal{T}$-invariant lines at $(0,0,k_{z})$ and $(\pi,\pi,k_{z})$, condition (III) reduces to $\lambda_{\rm so}\sin k_{z}=0$, which leads to $k_{z}=0$ and $\pi$, suggesting that only the four MKWNs appear on these two lines. On the line at $(0,\pi,k_{z})$, condition (III) simplifies to: $\lambda_{\rm so}\sin k_{z}+2\eta\lambda_{\rm M}=0$, which has solutions only if $\lambda_{\rm so}>2\eta\lambda_{\rm M}$. When this condition is fulfilled, two Weyl nodes emerge at $(0,\pi,-k_{0})$ and $(0,\pi,-\pi+k_{0})$, where $k_{0}=\arcsin 2\eta\lambda_{\rm M}/\lambda_{\rm so}$. Similarly, on the line at $(\pi,0,k_{z})$, two additional Weyl nodes emerge at $(\pi,0,k_{0})$ and $(\pi,0,\pi-k_{0})$, provided that $\lambda_{\rm so}>2\eta\lambda_{\rm M}$. A schematic of the these Weyl nodes and their associated Fermi arcs is provided in Fig. S1.
Figure S1: (a) and $\eta<\eta_{c}$ in (b). Topological charges and cone tilting details of the Weyl nodes are shown in the middle inset. (c-d) Solid (dotted) purple lines refer to Fermi arcs on the left (right) $x$-normal surface, and the black arrows denote their spin polarizations. The solid green rings represent the projections of bulk Fermi surface, with the chemical potential fixed at $\mu=0.2$. $\lambda_{\rm M}=0.4$ and $0.8$ in (c) and (d), respectively. Common parameters are $t=t_{z}=0$, $\lambda_{\rm so}=1.1$, and $\eta=1$.
@@
It is easy to see that besides the Weyl nodes discussed in the first scenario, additional Weyl nodes may emerge along the following lines:
@@
(1) On the line at $(k_{x},0,\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}})$: Condition (III) reduces to
@@
In this section, we present a specific magnetic configuration as an example to demonstrate a possible realization of the tight-binding Hamiltonian under consideration. The distribution of magnetic and nonmagnetic atoms is presented in Figure 1 in the main text, where the blue (white) spheres represent magnetic (nonmagnetic) atoms, and the red arrows indicate the orientations of local magnetic moments. We assume that the outer-shell electrons of magnetic atoms—responsible for forming local moments—are strongly localized, whereas those of nonmagnetic atoms are itinerant and govern the system’s transport properties. Because of this separation in localization and delocalization, our model describes the behavior of these itinerant electrons moving in the magnetic background formed by these magnetic atoms. When the hopping paths of these itinerant electrons intersect with the localized magnetic moments, they experience a spin-dependent potential. This potential, in turn, gives rise to spin-dependent hopping amplitudes as electrons with spins parallel or antiparallel to the localized magnetic moment exhibit different tunneling probabilities. For instance, the representative hoppings in the $xy$, yzyz, and $zx$ planes are shown in Figure 1(c-e) in the main text. In the $xy$ plane, hopping along the $x$ and $y$ directions becomes spin-polarized as $\pm\sigma_{z}$; while in the yzyz (xzxz) plane, diagonal hopping paths acquire $\pm\sigma_{y}$ ($\pm\sigma_{x}$) spin dependence. Key real-space symmetries of this structure includes three mirror symmetries $\mathcal{M}_{z}$, $\mathcal{M}_{xy}$, $\mathcal{M}_{\bar{x}y}$, two twofold roto-time-reversal symmetries $C_{2}\mathcal{T}$ about the principal axes ($x$, $y$), and a fourfold roto-time-reversal-symmetry $C_{4z}\mathcal{T}$. These symmetries dictate the allowed hoppings and band degeneracies of the system. Based on the magnetic configuration illustrated in Figure 1 in the main text, the tight-binding Hamiltonian describing the itinerant electrons (assuming a single orbital degree of freedom for these electrons) is given by
$H$ $=$ $-\sum_{\braket{i,j},\sigma}t_{ij}c_{i,\sigma}^{\dagger}c_{j,\sigma}+i\lambda_{\rm so}\sum_{\braket{i,j},\sigma,\sigma^{{}^{\prime}}}\boldsymbol{d}_{ij}\cdot c_{i,\sigma}^{\dagger}\boldsymbol{\sigma}_{\sigma\sigma^{{}^{\prime}}}c_{j,\sigma^{{}^{\prime}}}$ (S36)
$+\lambda_{\rm M}\eta\sum_{\braket{i,j},\sigma\sigma^{{}^{\prime}}}\boldsymbol{S}_{ij}\cdot c_{i,\sigma}^{\dagger}\boldsymbol{\sigma}_{\sigma\sigma^{{}^{\prime}}}c_{j,\sigma^{{}^{\prime}}}+\lambda_{\rm M}\sum_{\braket{\braket{i,j}},\sigma\sigma^{{}^{\prime}}}\boldsymbol{S}_{ij}\cdot c_{i,\sigma}^{\dagger}\boldsymbol{\sigma}_{\sigma\sigma^{{}^{\prime}}}c_{j,\sigma^{{}^{\prime}}}+h.c.$
$=$ $\sum_{\boldsymbol{k}}\Psi_{\boldsymbol{k}}^{\dagger}\mathcal{H}(\boldsymbol{k})\Psi_{\boldsymbol{k}}.$
@@
Here, $c_{i,\sigma}(c_{i,\sigma}^{\dagger})$ represents the annihilation (creation) operator for an electron with spin $\sigma$ at site $i$. The notation $\braket{i,j}$ indicates nearest-neighbor hopping between sites $i$ and $j$, $\braket{\braket{i,j}}$ indicates next-nearest-neighbor hopping, and the unit vector $\boldsymbol{d}_{ij}$ points along the bond direction from site $j$ to site $i$. The parameter $t_{ij}$ refers to the hopping amplitude between two nearest-neighbor sites, and $\lambda_{\rm M}$ characterizes the difference in hopping amplitude for opposite spins that is induced by the background magnetic moments. $\boldsymbol{S}_{ij}$ denotes the local magnetic moment along the hopping path between site $i$ and $j$. The spin-orbit coupled terms of the amplitude $\lambda_{\rm so}$ arise when there are perturbations that break the inversion and three mirror symmetries. By performing a Fourier transformation to the momentum space and choosing the basis as $\Psi_{\boldsymbol{k}}^{\dagger}=(c_{\uparrow,\boldsymbol{k}}^{\dagger},c_{\downarrow,\boldsymbol{k}}^{\dagger})$, we obtain the momentum-space Hamiltonian, which reads
$\mathcal{H}(\boldsymbol{k})$ $=$ $-\left[t(\cos k_{x}+\cos k_{y})+t_{z}\cos k_{z}\right]+\lambda_{\rm so}(\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y}+\sin k_{z}\sigma_{z})$ (S37)
$+\lambda_{\rm M}\eta(\cos k_{x}-\cos k_{y})\sigma_{z}+\lambda_{\rm M}\sin k_{z}(-\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y}),$
where the Pauli matrices $(\sigma_{x},\sigma_{y},\sigma_{z})$ acts on the spin degrees of freedom. Importantly, our model Eq. (S36) includes all symmetry-allowed nontrivial terms from nearest- and next-nearest-neighbor hoppings. Longer-range spin-independent contributions—such as $\sin k_{z}(\sin k_{x}+\sin k_{y})\sigma_{0}$ terms—are omitted because they do not affect the system’s topology. Those longer-range spin-dependent terms—such as $(\cos 2k_{x}-\cos 2k_{y})\sigma_{z}$—are also neglected since their amplitudes are expected to be negligible within the tight-binding framework and do not qualitatively alter the results.
Weyl node position Charge Requirements for existence
$\left(0,0,-\arcsin\frac{B_{z}}{\lambda_{\rm so}}\right)$ +1+1 $|B_{z}|<\lambda_{\rm so}$
$\left(0,0,-\pi+\arcsin\frac{B_{z}}{\lambda_{\rm so}}\right)$ 1-1
$\left(\pi,\pi,-\arcsin\frac{B_{z}}{\lambda_{\rm so}}\right)$ +1+1
$\left(\pi,\pi,-\pi+\arcsin\frac{B_{z}}{\lambda_{\rm so}}\right)$ 1-1
$\left(0,\pi,-\arcsin\frac{B_{z}+2\lambda_{\rm M}\eta}{\lambda_{\rm so}}\right)$ 1-1 $\eta<\eta_{c}-\frac{B_{z}}{2\lambda_{\rm M}}$
$\left(0,\pi,-\pi+\arcsin\frac{B_{z}+2\lambda_{\rm M}\eta}{\lambda_{\rm so}}\right)$ +1+1
$\left(\pi,0,\arcsin\frac{2\lambda_{\rm M}\eta-B_{z}}{\lambda_{\rm so}}\right)$ 1-1 $\eta<\eta_{c}+\frac{B_{z}}{2\lambda_{\rm M}}$
$\left(\pi,0,\pi-\arcsin\frac{2\lambda_{\rm M}\eta-B_{z}}{\lambda_{\rm so}}\right)$ +1+1
$\left(\pm\arccos(1-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}-\frac{B_{z}}{\lambda_{\rm M}\eta}),0,\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ 1-1 $\lambda_{\rm so}<\lambda_{\rm M}$, $\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}}<B_{z}<\lambda_{\rm M}(2\eta-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}})$
$\left(\pm\arccos(1-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}-\frac{B_{z}}{\lambda_{\rm M}\eta}),0,\pi-\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ +1+1
$\left(\pm\arccos(1+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}+\frac{B_{z}}{\lambda_{\rm M}\eta}),\pi,\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ 1-1 $\lambda_{\rm so}<\lambda_{\rm M}$, $\lambda_{\rm M}(2\eta+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}})<B_{z}<-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}}$
$\left(\pm\arccos(1+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}+\frac{B_{z}}{\lambda_{\rm M}\eta}),\pi,\pi-\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ +1+1
$\left(\pi,\pm\arccos(1+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}-\frac{B_{z}}{\lambda_{\rm M}\eta}),-\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ 1-1 $\lambda_{\rm so}<\lambda_{\rm M}$, $\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}}<B_{z}<\lambda_{\rm M}(2\eta+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}})$
$\left(\pi,\pm\arccos(1+\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}-\frac{B_{z}}{\lambda_{\rm M}\eta}),-\pi+\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ +1+1
$\left(0,\pm\arccos(1-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}+\frac{B_{z}}{\lambda_{\rm M}\eta}),-\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ 1-1 $\lambda_{\rm so}<\lambda_{\rm M}$, $\lambda_{\rm M}(2\eta-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}})<B_{z}<\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}}$
$\left(0,\pm\arccos(1-\frac{\lambda_{\rm so}^{2}}{\lambda_{\rm M}^{2}\eta}+\frac{B_{z}}{\lambda_{\rm M}\eta}),-\pi+\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}}\right)$ +1+1
Table 1: Positions, charges, and existence conditions for Weyl nodes under a Zeeman field in the $z$ direction. $\eta_{c}=\lambda_{\rm so}/2\lambda_{\rm M}$.
@@
In the main text, we have shown that, when $\lambda_{\rm so}=0$, a weak Zeeman field will change the CNLs to two nodal rings located at the two mirror-invariant planes at $k_{z}=0$ and $\pi$, leading to the occurrence of three-dimensional quantized anomalous Hall effects when the Fermi surface coincides with these two nodal rings. The underlying physics can be understood by the layer Chern number. Specifically, from a dimension-reduction perspective, the two-dimensional layer Hamiltonian with a fixed $k_{z}$ describes a Chern insulator characterized by a Chern number of value $\mathcal{C}=1$ or 1-1 (depending on the direction of the Zeeman field) when $k_{z}\neq 0$ and $\pi$. Accordingly, each layer contributes to a Hall conductivity $\sigma_{xy}=\frac{e^{2}}{h}$. By summing the contribution from all layers, one obtains